OpenStudy (anonymous):

-(2x - 3y)+ 5x + 2 - 3x + 3y + 2 simplify

OpenStudy (anonymous):

4 + 6y \(\huge\cal\rlap{\color{lime}-\color{lime}G\color{lime}r\color{lime}e\color{lime}e\color{lime }n\color{lime}ツ}{\color{black}{\;-Green~ツ}}\)

OpenStudy (anonymous):

Welcome to Open Study! You can give medals by clicking 'Best Response'.

OpenStudy (anonymous):

first distribute thr negative through the parenthesis -2x + 3y + 5x + 2 - 3x + 3y + 2 now just combine all your like terms -2x - 3x + 5x = 0 (they cancel) 3y + 3y = 6y 2 + 2 = 4 now put it all together 6y + 4